OpenStudy (anonymous):

Which graph represents the solutions to the inequality |2x - 4| greater than or equal to 4?

OpenStudy (solomonzelman):

I am unable to access the attachment -:( BUT.... \(\Large\color{blue}{ \bf |2x - 4|=4}\) \(\Large\color{blue}{ \bf 2x - 4=4~~~~~~~or~~~~~~~2x-4=-4}\) \(\Large\color{blue}{ \bf 2x =8~~~~~~~~~~~~~~~~~~~~~~~~~2x=0}\) \(\Large\color{blue}{ \bf x =4~~~~~~~~~~~~~~~~~~~~~~~~~~~x=0}\)